OpenStudy (anonymous):

hELP I REALLY NEED HELP Add the following fractions -3 1/2+4 2/3

OpenStudy (anonymous):

Hi

OpenStudy (anonymous):

Let's see what you've learned. What would be a common denominator or 2 and 3?

OpenStudy (anonymous):

i have two more then im done i really need help

OpenStudy (anonymous):

*of, not or

OpenStudy (anonymous):

6

OpenStudy (anonymous):

Yes, so can you make then new fractions with a denominator of 6?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Cool, give it a shot. :)

OpenStudy (anonymous):

-9/6 + 16/6

OpenStudy (anonymous):

7/6

OpenStudy (anonymous):

or 1 1/6

OpenStudy (anonymous):

We're not going to worry about the whole numbers just yet, Just the fractions. So, they have a common denominator of 6. 6/2=3, so multiply the numerator by 3 1*3=3 So, the new fraction is 3/6 6/3=2, so multiply the numerator by 2 2*2=4 So, the new fraction is 4/6 Understand, or do you need more detail?

OpenStudy (anonymous):

oh ok i get it is that the answer

OpenStudy (anonymous):

No, that is just the equivalent fractions that will help us to get the answer. So, our new problem is:\[-3\frac{ 3 }{ 6 }+4\frac{ 4 }{ 6 }\]Think you can solve it?

OpenStudy (md152727):

so -3+4=1 right?

OpenStudy (anonymous):

yes

OpenStudy (md152727):

and 3+4?

OpenStudy (anonymous):

correct

OpenStudy (anonymous):

Then what is your overall answer?

OpenStudy (md152727):

1 and 3+4

OpenStudy (md152727):

1 and 3+4/6

OpenStudy (md152727):

just need to add her answer

OpenStudy (anonymous):

That is not correct.

OpenStudy (md152727):

@Mc2 what is 3+4

OpenStudy (anonymous):

-15/6+28/6

OpenStudy (anonymous):

@MD152727, please stop trying to lead @Mc2 to the incorrect answer.

OpenStudy (md152727):

I know the answer already

OpenStudy (anonymous):

13/6

OpenStudy (anonymous):

is this correct

OpenStudy (md152727):

nope

OpenStudy (anonymous):

\[-3\frac{ 3 }{ 6 }+4\frac{ 4 }{ 6 }\]Let's start with the whole numbers first. \[-3+4=1\]Next is the fractions. To add them, simply do so with the numerators.\[-3+4=1\]This means that our fractions add to:\[\frac{ 1 }{ 6 }\]Add our whole number to the fraction to get the answer:\[1+\frac{ 1 }{ 6 }=1\frac{ 1 }{ 6 }\]

OpenStudy (md152727):

@mc2 it is so easy!

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

Subtract the two following Fractions -2 3/8-(-4 1/2)

OpenStudy (md152727):

you should know how to solve these now

OpenStudy (anonymous):

@Mc2 Do you want to give it a shot, and then I'll check your work? :)

OpenStudy (anonymous):

sure i have to turn this in at 5:00 though

OpenStudy (anonymous):

O.o It is 7:47 my time. What time is it where you are?

OpenStudy (md152727):

yeah, 7:48 pm my time

OpenStudy (anonymous):

4:48

OpenStudy (md152727):

OMG, how many questions?

OpenStudy (anonymous):

Well, better get solvin'

OpenStudy (anonymous):

-19/8+12/8

OpenStudy (anonymous):

Sorta...

OpenStudy (anonymous):

i mean -19/8-(-12/8)

OpenStudy (anonymous):

-19/8 is correct, but 2*4=8 8+1=9 So, the second should be -9/8

OpenStudy (md152727):

simplified: \[-19+12/8\]

OpenStudy (md152727):

add -19+12

OpenStudy (md152727):

what do you get?

OpenStudy (md152727):

@Mc2

OpenStudy (anonymous):

-19/8-(-9/8)=-10/8

OpenStudy (md152727):

nope

OpenStudy (md152727):

you should get -7/8

OpenStudy (md152727):

Ok, 6 minutes left, how many questions?

OpenStudy (anonymous):

\[-\frac{ 19 }{ 8 }-(-\frac{ 9 }{ 8 })\]\[-19--9=-10\]\[-\frac{ 10 }{ 8 }=-1\frac{ 1 }{ 4 }\]

OpenStudy (md152727):

didn't she say -19/8-(-12/8)

OpenStudy (md152727):

are you going to make me crazy??

OpenStudy (md152727):

i mean -19/8-(-12/8)

OpenStudy (md152727):

that is what she said

OpenStudy (anonymous):

i sent u a message orange master

OpenStudy (md152727):

\(\color{blue}{\text{Originally Posted by}}\) @Mc2 i mean -19/8-(-12/8) \(\color{blue}{\text{End of Quote}}\) WHAT IS THE QUESTION

OpenStudy (anonymous):

Subtract the two following Fractions -2 3/8-(-4 1/2)

OpenStudy (anonymous):

thanks guys bye

OpenStudy (anonymous):

Do you see @MD152727

OpenStudy (anonymous):

will u be on tommorow

OpenStudy (anonymous):

Prob not

OpenStudy (md152727):

maybe

OpenStudy (anonymous):

oh ok

OpenStudy (md152727):

but tag us

OpenStudy (md152727):

bye!

OpenStudy (anonymous):

@OrangeMaster

OpenStudy (md152727):

??why are you tagging yoursellf

OpenStudy (anonymous):

If I'm on, but prob not :( To show them in case they didn't know

OpenStudy (md152727):

oh ok

OpenStudy (anonymous):

Bye, got to go to work in morning, so I'm going to sleep.